The work required to compress a gas reversibly according to pV^1.30 = c is 67,790J, if there is no flow. Determine ΔU and Q if the gas is (a) air, (b) methane. For methane, k= 1.321, R=518.45 J/kg., Csubv= 1.6187, csubp= 2.1377 kJ/kg.K.
1
Expert's answer
2020-10-25T18:28:53-0400

(a) The amount of heat transferred by the air is,

"Q=(\frac{K_{air} - n}{K_{air} - 1})W"

"K_{air}=" specific heat ratio of air

"W=" work required to compress the gas reversibly

substitute 1.4 for "K_{air}"", 1.3" for "n" and "-67,790J" for "W"

"Q= (\frac{1.4-1.30}{1.4-1})(-67,790J)=-16947.5J=-16.95kJ"

The amount of heat transferred by the air ("Q" ) is "16.95kJ"


change in internal energy of the air is,

"Q=\Delta U+W"

"Q=-16947.5J" and "W=-67,790J"

"-16947.5J=\Delta U-67,790"

"\Delta U=-16947.5J+67,790J"

"\Delta U = 50842.5J= 50.84kJ"


(b) Amount of heat transferred by the methane,

"Q=(\frac{K_{methane}-n}{K_{methane}-1})W"

"K_{methane}=1.321, n=1.30 ,W=-16,790J"

"Q=(\frac{1.321-1.30}{1.321-1})(-67,790)=-4434.86J=-4.43kJ"


the amount of heat transferred by the methane ("Q" ) is "4.43kJ"


Change in internal energy of the methane,

"Q=\Delta U+W"

"-4434.86J=\Delta U-67,790J"

"\Delta U= -4434.86J+67,790J=63355.14J=63.36kJ"